[Brahms-l] Shift report 20050610 16:00-24:00

From: <apache@brahms-web.brahms.bnl.gov>
Date: Fri Jun 10 2005 - 00:16:03 EDT
Shifter leader was rk
Shifters were 


1. Beam from previous shift lost at 19.27 due to quench link interlock, another injection at 22.00, beam still in RHIC.
2. Moved FFS to 3 deg and BFS to 2.3, magnets are all at full field, polarity A. 
Runs:
15342 - 15347 (FFS @ 30 deg, 1/8 B; MRS @ 34 deg, 4kGA) 511k tr.3, 19k tr.6
15352, 15353 (FFS @ 3, BFS @ 2.3, MRS @ 34, all full field A) 50k tr.2 206k tr.3
3. FFS1440 mainframe had some problems - resetted, T3_A2xu tripped once.
No other problems seen.

_______________________________________________
Brahms-l mailing list
Brahms-l@lists.bnl.gov
http://lists.bnl.gov/mailman/listinfo/brahms-l
Received on Fri Jun 10 00:16:16 2005

This archive was generated by hypermail 2.1.8 : Fri Jun 10 2005 - 00:16:27 EDT